# Problem-informed Graphical Quantum Generative Learning

### v1

@@ V Acknowledgements
Figure 8: Decomposition of $U_{Z}$ operators generated by many-body Ising Hamiltonians. (a) Complete $U_{Z}(\boldsymbol{\alpha})$ operator corresponding to MN in Fig. 2b. (b) Decomposition of $\exp{(-i\alpha ZZZ)}$, a $3$-body $MultiRZ$ gate parametrized by $\alpha$. (c) Implementation of all terms of a clique of size $3$ with an additional ancillary qubit. In all the figures, parametrized gates represent $k$-local $MultiRZ$ operators.
Appendix A Implementation details
@@
A.1 Quantum Circuit Markov Random Fields
Markov networks define higher-order Ising Hamiltonians as described in III.1, which generate QCMRF circuits composed of $MultiRZ$ gates. For the MN in Fig. 2b, the corresponding parametrized circuit is shown in Fig. 8a. A $k$-local $MultiRZ$ gate can be implemented in linear depth with $2k$ $CNOT$ gates and a single-qubit $R_{Z}$ rotation. An example for this decomposition in shown in Fig. 8b. There are several other alternative strategies for implementing these circuits, as explained in the context of QAOA in Refs. [57, 58]. For example, one could implement the circuit corresponding to a clique with $m$ random variables with significantly less $CNOT$ gates using an ancillary qubit, as shown in Fig. 8c. This approach reduces the number of gates and the depth, while adding an ancillary qubit for each clique of size $>2$.
@@ A.1 Quantum Circuit Markov Random Fields
A.2 Bayesian Quantum Circuits
Here we start by describing the general idea of BQCs introduced in [33]. This model associates a qubit to each binary random variable of the BN and then applies unitary operations in the following manner. First parametrized single-qubit $R_{Y}$ rotation gates are applied to qubits for which the corresponding nodes have no parents. Then uniformly controlled $R_{Y}$ operations are performed on each qubit, where the control qubits correspond to the parents of the given node. Note, that since a node can have multiple parents, this can lead to gates with a large number of controls. The order of application of these unitaries has to follow a two rules: every qubit can only be targeted once and after a qubit was used as a control, it cannot be targeted anymore. These rules ensure the compliance with the directed acyclic nature of the graph. The basis-enhancement introduced in [35] further applies parametrized general $U_{f}(\Gamma,\Delta,\Sigma)$ gates to all qubits. The BBQC corresponding to the BN presented in 2a is shown in Fig. 9a.
While uniformly controlled gates provide an easily interpretable mapping from the BN graph to a parametrized quantum circuit, they cannot be implemented directly. The number of parameters for such a gate is $2^{n_{c}}$, where $n_{c}$ is the number of control qubits. In [34] and [46], the authors gave several strategies to decompose these operators. For the purpose of this work, it is easiest to think of these as the decomposition in terms of (multi-controlled) $R_{Y}$ rotations and $X$ gates. This decomposition of a uniformly controlled $R_{Y}$ rotation with $2$ controls is shown in Fig. 9b. The multi-controlled gates can be further decomposed into single qubit rotation and $CNOT$ gates with additional ancillary qubits as described in [34] or by adapting the ancilla-free strategy of [59].
@@ A.2 Bayesian Quantum Circuits
Appendix B Resource estimation
@@ B.1 Number of parameters
In the case of the QCIBM and QCMRF models, the number of parameters depends mostly on the number of terms in Hamiltonian that generates the time evolution. However, these models cannot be compared directly, since in QCIBMs this number only depends on the number of qubits $n$ and in QCMRF circuits there is an explicit dependence on the graph topology and the clique factorization, but also the overlap of the cliques.
@@
It is worth mentioning, that for pairwise MNs, meaning those defined by only pairwise interactions lead to two-body terms between the qubits representing connected nodes and one-body Pauli-Z terms on all qubits. These give rise to QCMRF circuits with less or equal number of parameters than the corresponding QCIBM Ansatz, where we have equality only for complete graphs.
@@ B.2 Number of qubits
B.3 Circuit depth
Based on the implementation of multi-qubit gates, the depth of the circuit can also vary. In general, the more ancillas we use for decomposing these unitaries, the shallower the final circuit can be. For a fair comparison, here we only consider ancilla-free implementations and analyse the depth accordingly. We also assume full connectivity of qubits and parallel execution of gates that act on disjoint sets of qubits, where possible and use parametrized one-qubit $U_{f}$ and $CNOT$ gates as our basis gate set.
@@ B.3 Circuit depth
For QCMRF circuits, the scaling is more complicated, since it depends on the clique-structure of the MN. Here we give an upper-bound on the depth, considering the Hamiltonian before reparametrization. For each clique $C$, we have $12^{|C|}-$ 1 $MultiRZ$ gates, and there are (${|C|\choose k}$ $k$-local gates for each $k\leq|C|$. To minimize the depth, we can implement two gates in parallel in each step, such that the qubits they act on are complements of each other in the clique. In the ancilla-free case, each $k$-local gate can be implemented in $12k-$ 1 depth. This means, that the depth required to implement such a clique is $\mathcal{O}(2^{|C|})$. If we assume $|\mathcal{C}|$ cliques in the graph and the size of the largest maximal clique is $l_{m}$, than the depth required is $\mathcal{O}(|\mathcal{C}|2^{l_{m}})$. However, this is a crude upper bound in the worst case, not taking into account the parallel execution of gates of disjoint cliques. In practice, this scaling can be much better.
@@
We compare these two models on the simple yet illustrative example of a graph composed of a single triangle. This is obviously a chordal graph, and the number of trainable parameters is equal in both circuit. In this case, the depth of the QCMRF circuit is $16$. In the corresponding BN, the first node has no parents leading to depth $1$. The second node has $1$ parent, leading to $2$ single-controlled $R_{Y}$ gates, each implemented in depth $4$, thus depth $10$ with the $X$ gates. The final node has $2$ parents, meaning $4$ multi-controlled gates with two controls, each having depth $14$, thus depth $60$ in total. The final depth of this circuit is $72$ with the $U_{f}$ gates before measurement. This example shows that BBQCs can be much deeper than QCMRF circuits on chordal graphs, while this is not reflected in the number of trainable parameters.
@@
$\begin{split}P(X_{1},\dots,X_{n})&=P(X_{1})P(X_{2}|X_{1})P(X_{3}|X_{1},X_{2})\dots\\ &\cdot P(X_{n}|X_{n-k+1},\dots,X_{n-1}).\end{split}$ (13)

### v2

@@ Appendix B Implementation details
Figure 7: Decomposition of $U_{Z}$ operators generated by many-body Ising Hamiltonians. (a) Complete $U_{Z}(\boldsymbol{\alpha})$ operator corresponding to MN in Fig. 2b. (b) Decomposition of $\exp{(-i\alpha ZZZ)}$ ), a $3$-body $MultiRZ$ gate parametrized by $\alpha$. (c) Implementation of all terms of a clique of size $3$ with an additional ancillary qubit. In all the figures, parametrized gates represent $k$-local $MultiRZ$ operators.
@@
B.1 Quantum Circuit Markov Random Fields
Markov networks define higher-order Ising Hamiltonians as described in III.1, which generate QCMRF circuits composed of $MultiRZ$ gates. For the MN in Fig. 2b, the corresponding parametrized circuit is shown in Fig. 7a. A $k$-local $MultiRZ$ gate can be implemented in linear depth with $2k$ $CNOT$ gates and a single-qubit $R_{Z}$ rotation. An example for this decomposition in shown in Fig. 7b. There are several other alternative strategies for implementing these circuits, as explained in the context of QAOA in Refs. [61, 62]. For example, one could implement the circuit corresponding to a clique with $m$ random variables with significantly fewer $CNOT$ gates using an ancillary qubit, as shown in Fig. 7c. This approach reduces the number of gates and the depth, while adding an ancillary qubit for each clique of size $>2$.
@@ B.1 Quantum Circuit Markov Random Fields
B.2 Bayesian Quantum Circuits
Here we start by describing the general idea of BQCs introduced in Ref. [34]. This model associates a qubit to each binary random variable of the BN and then applies unitary operations in the following manner. First parametrized single-qubit $R_{Y}$ rotation gates are applied to qubits for which the corresponding nodes have no parents. Then uniformly controlled $R_{Y}$ operations are performed on each qubit, where the control qubits correspond to the parents of the given node. Note, that since a node can have multiple parents, this can lead to gates with a large number of controls. The order of application of these unitaries has to follow two rules: every qubit can only be targeted once and after a qubit was used as a control, it cannot be targeted any more. These rules ensure the compliance with the directed acyclic nature of the graph. The basis-enhancement introduced in Ref. [36] further applies parametrized general $U_{f}(\Gamma,\Delta,\Sigma)$ gates to all qubits. The BBQC corresponding to the BN presented in 2a is shown in Fig. 8a.
While uniformly controlled gates provide an easily interpretable mapping from the BN graph to a parametrized quantum circuit, they cannot be implemented directly. The number of parameters for such a gate is $2^{n_{c}}$, where $n_{c}$ is the number of control qubits. In Ref. [35] and [47], the authors gave several strategies to decompose these operators. For the purpose of this work, it is easiest to think of these as the decomposition in terms of (multi-controlled) $R_{Y}$ rotations and $X$ gates. This decomposition of a uniformly controlled $R_{Y}$ rotation with $2$ controls is shown in Fig. 8b. The multi-controlled gates can be further decomposed into single qubit rotation and $CNOT$ gates with additional ancillary qubits as described in Ref. [35] or by adapting the ancilla-free strategy of [63].
@@ B.2 Bayesian Quantum Circuits
Appendix C Resource estimation
@@ C.1 Number of parameters
In the case of the QCIBM and QCMRF models, the number of parameters depends mostly on the number of terms in the Hamiltonian that generates the time evolution. However, these models cannot be compared directly, since in QCIBMs this number only depends on the number of qubits $n$ and in QCMRF circuits there is an explicit dependence on the graph topology and the clique factorization, but also the overlap of the cliques.
@@
It is worth mentioning, that for pairwise MNs, meaning those defined by only pairwise interactions, lead to two-body terms between the qubits representing connected nodes and one-body Pauli-Z terms on all qubits. These give rise to QCMRF circuits with less or equal number of parameters than the corresponding QCIBM Ansatz, where we have equality only for complete graphs.
@@ C.2 Number of qubits
C.3 Circuit depth
Based on the implementation of multi-qubit gates, the depth of the circuit can also vary. In general, the more ancillas we use for decomposing these unitaries, the shallower the final circuit can be. For a fair comparison, here we only consider ancilla-free implementations and analyse the depth accordingly. We also assume full connectivity of qubits and parallel execution of gates that act on disjoint sets of qubits, where possible, and use parametrized one-qubit $U_{f}$ and $CNOT$ gates as our basis gate set.
@@ C.3 Circuit depth
For QCMRF circuits, the scaling is more complicated, since it depends on the clique-structure of the MN. Here we give an upper-bound on the depth, considering the Hamiltonian before reparametrization. For each clique $C$, we have $12^{|C|}-$ 1 $MultiRZ$ gates, and there are (${|C|\choose k}$ $k$-local gates for each $k\leq|C|$. To minimize the depth, we can implement two gates in parallel in each step, such that the qubits they act on are complements of each other in the clique. In the ancilla-free case, each $k$-local gate can be implemented in $12k-$ 1 depth. This means, that the depth required to implement such a clique is $\mathcal{O}(2^{|C|})$. If we assume $|\mathcal{C}|$ cliques in the graph and the size of the largest maximal clique is $l_{m}$, then the depth required is $\mathcal{O}(|\mathcal{C}|2^{l_{m}})$. However, this is a crude upper bound in the worst case, not taking into account the parallel execution of gates of disjoint cliques. In practice, this scaling can be much better.
@@
We compare these two models on the simple, yet illustrative example of a graph composed of a single triangle. This is obviously a chordal graph, and the number of trainable parameters is equal in both circuit. In this case, the depth of the QCMRF circuit is $16$. In the corresponding BN, the first node has no parents leading to depth $1$. The second node has $1$ parent, leading to $2$ single-controlled $R_{Y}$ gates, each implemented in depth $4$, thus depth $10$ with the $X$ gates. The final node has $2$ parents, meaning $4$ multi-controlled gates with two controls, each having depth $14$, thus depth $60$ in total. The final depth of this circuit is $72$ with the $U_{f}$ gates before measurement. This example shows that BBQCs can be much deeper than QCMRF circuits on chordal graphs, while this is not reflected in the number of trainable parameters.
Figure 10: QCMRF benchamrk results against the problem agnostic QCIBM model using KL divergence and exact quantum circuit simulation. These training curves show the evolution of the cost function leading to the final TV distance in Fig. 11.
@@
$\begin{split}P(X_{1},\dots,X_{n})&=P(X_{1})P(X_{2}|X_{1})P(X_{3}|X_{1},X_{2})\dots\\ &\cdot P(X_{n}|X_{n-k+1},\dots,X_{n-1}).\end{split}]$ (17)
@@
Appendix D Additional numerical simulations
D.1 Scaling with exact simulation
We also present experiments based on exact simulation, i.e., exact probabilities corresponding to infinitely many shots. Here we only consider the KL divergence based on the exact model and target probability distributions. While this training is not realistic, as it requires the exact probabilities from both distributions, this way, the simulation can be scaled up to larger problems, than the ones considered in the main text.
For this comparison, we use the triangle chain graph, that is equivalent to the $3$-gram model, as in Eq. (17), each clique being of size $3$. In Fig. 10, we present all the training curves with similar hyperparameters as in the simulation in the main text. We show the average of $5$ independent problems, having different sets of factor values. The performance scaling with the number of qubits is shown in Fig. 11. Here, we define the final TV distance as the average of the last $100$ training epochs in each case. In this figure, markers show the average final TV distance among the $5$ problems, and errorbars show the standard deviation around the mean.
Figure 11: Scaling of the final TV distance with the number of qubits using the KL divergence. Here we use triangle chain graphs, that correspond to $3$-gram models. The markers show the average final TV distance of $5$ independent problems and errorbars show the standard deviation.
The separation in the average performance between the QCMRF and QCIBM models is independent of the problem size, while the number of parameters is $\mathcal{O}(n)$ and $\mathcal{O}(n^{2})$ respectively. For $20$ qubits, this amounts to $135$ and $270$ parameters.
D.2 Comparison with classical methods
While we concentrate on comparing QCMRF to other quantum models, here we also provide a preliminary comparison with a classical method. We consider the grid graph topology with $9$ nodes, and turn the corresponding MN into a BN. The parameters of the resulting BN are fitted to the training set using maximum likelihood estimation, as implemented in the pgmpy Python package [66]. After fitting, the classical model to the data, we calculate the corresponding probability distribution exactly and compute the TV distance from the target distribution. We repeat this procedure for $100$, $500$ and $1000$ training samples. The corresponding results are shown in Fig. 12. Similarly, the QCMRF models are trained with the same number of training samples, using the MMD loss function. In each case, we run $5$ independent problems and plot the results similarly to previous experiments.
While our model shows similar performance with small training sets, its performance is weaker when the number of samples is increased. This highlights the need for more enhanced training methods for quantum models, however, this does not affect the potential of our model to have quantum advantage in sampling the learned probability distribution, since these classical methods are known to be unscalable.
Figure 12: Benchmark results against classical BNs. We consider training sets of size $100$, $500$ and $1000$ for both the classical and quantum models. The QCMRF models are trained with the MMD loss function. The TV distance is computed exactly in each case.